Solve the quadratic equation by completing the square:
x^2 - 6x - 3 = 0
Can someone give me a heads up on how to solve this please?

Answer by Nate(3500)

x^2 - 6x - 3 = 0
x^2 - 6x = 3
(x - 3)^2 = 3 + 9 = 12
x - 3 = +- sqrt(12)
x = 3 +- 2*sqrt(3)